# Real-World Data and Calibrated Simulation Suite for Offline Training of Reinforcement Learning Agents to Optimize Energy and Emission in Buildings for Environmental Sustainability

### v1

@@ 8 Demonstration Benchmarking Results
Training a Reinforcement Learning Agent on the Simulator To demonstrate the usefulness of our calibrated simulator on generating an improved policy, we used Soft Actor Critic (SAC) algorithm (Haarnoja et al., 2018) to train an agent, and then compared our agent with the baseline performance of running the policy currently used in the real building. Both actor and critic were feedforward networks. We ran hyperparameter tuning, again using the method from Golovin et. al. (Golovin et al., 2017), to choose the dimensionality of the critic network and actor network, the batch size, the critic learning rate and actor learning rate, and $\gamma$.
@@
We recorded the actor loss, critic loss, alpha loss, and return, over a two day period. The agents trained for 4,000 iterations. Using the $R_{3C}$ reward, the baseline over this two day period had a return of -12.9, and our best agent had an improved return of -11.9, an 8% improvement over the baseline, as show in Table 3. For further training details, and an in depth performance comparison between the learned policy and the baseline, including a breakdown on setpoint deviation, carbon emissions, electrical energy, and natural gas energy, see Appendix H.
Training a Learned Dynamics Model Another important task is to use a sequence model to learn to predict the real world data, effectively learning a dynamics model that can then be used in turn in place of the simulator to train an agent. To demonstrate this approach, we trained an encoder-decoder LSTM(Hochreiter, 1997) to model the building dynamics. The model takes in a historical sequence of length $N$ and outputs a prediction sequence of length $M$. At each timestep $t$ in the sequence, the model is given an observation $O_{t}$, action taken by the policy $A_{t}$, and auxiliary state features (such as time of day and weather, that are useful as inputs but need not be predicted) $U_{t}$, and for future timesteps, the model is trained to predict future observations, as well as future reward information (based on predicted energy use and carbon emissions) $E_{t}$. We evaluated this model by comparing its predictions with the real world data over a three week period, finding that it achieved strong performance and successfully modeled many building dynamics. For detailed architecture diagrams, training information and performance analysis, see Appendix I.
Training a Reinforcement Learning Agent on Real Data Building directly off of the above, we also trained an RL agent on the learned dynamics model, demonstrating the ability to learn a policy directly from data without involving the simulator. Like the simulator SAC agent, we were able to learn a policy that improved upon the baseline. For detailed analysis of this policy, see Appendix J.
@@ 9 Limitations and Conclusion
The biggest limitation of our benchmark is that all buildings are located in California. We intend to remedy this in the near future by adding more buildings. Another limitation is that we only include data from a one year duration, and in the future we may add longer sequences, for year over year analysis. Our simulator also lacks a radiative heat model, and we hope further work can add this. In addition, our calibration focused on temperature, but in future work we hope to include energy consumption metrics as part of the calibration procedure.
@@
We present a high quality interactive HVAC Control Suite, with real-world historical data from three buildings, as well as calibrated simulators for each building, and a novel, data-based, simulation calibration procedure. We also show promising initial results on key benchmark tasks. We believe this benchmark will facilitate collaboration, reproducibility, and progress on this problem, making an important contribution towards environmental sustainability.
Appendix A Reward Function Details
@@
We call our reward function the 3C Reward, because it is made up of a combination of three facors: Comfort, Cost, and Carbon. The purpose of the reward function is to provide the agent a feedback signal after each action about the quality of the current and past actions performed. We combine the different objectives described in Optimization Problem as a normalized, weighted sum of maintaining comfort conditions, electrical cost, and carbon cost:
@@
where $C_{1}$ represents normalized comfort conditions, $C_{2}$ normalized energy cost and $C_{3}$ normalized carbon emission. Constants $u$, $v$, $w$ represent operator preferences, allowing them to weight the relative importance of cost, comfort and carbon consumption.
Each value $C_{1},C_{2},C_{3}$, is bounded by the range $[-1,0]$, where worst performance is $-1$ and the ideal performance upper-bound is 00 Thus the reward function in an agregate is formulated as an approximate regret function, bounded in the range [-1,0], and represents an offset from the best-case where comfort conditions are perfectly maintained, without consuming energy and emitting carbon. Each of the sub functions $C_{1},C_{2},C_{3}$ will be elaborated next.
A.1 Comfort Loss Function ($C_{1}$)
@@
Studies have shown that a relationship exists between work performance and temperature. For example, in Seppänen, et al. 2006 (Seppanen et al., 2006), work performance was quantified as the mean time required to complete common office tasks (e.g., text processing, bookkeeping calculations, telephone customer service calls, etc.). Performance was shown to increase gradually with temperatures increasing up to 21-22℃ and decreasing at temperatures beyond 23-24℃. Therefore, when temperatures deviate outside setpoints, the comfort loss should also be smooth and monotonically increasing.
@@
The comfort loss function represents a bounded penalty term for occupied zones that have zone air temperatures outside of setpoint, and covers three adjacent temperature intervals: below cooling setpoint $T_{z}<\hat{T}_{heating}$, inside setpoints $\hat{T}_{heating}\leq T_{z}\leq\hat{T}_{cooling}$, and above cooling setpoint $\hat{T}_{cooling}<T_{z}$
@@
Live Occupant Feedback The idea of human feedback shaping the agent’s policy may be particularly suitable for the smart buildings project, and has been detailed in Knox and Stone 2009. While not implemented in the initial version of the reward function, the comfort loss function can be extended with an occupant feedback signal reflecting discomfort (e.g., “too hot” or “too cold”) in a variety of methods like Mozer 1998 (Mozer, 1998). The agent’s goal should be to minimize this type of feedback, and the regret should be increased anytime this feedback signal is received. Suppose one or more occupants in zone $z$, provided a “too cold” feedback signal, $\hat{T}_{heating}$ may be increased by a small amount from the baseline setpoint configuration, and may smoothly return to the baseline smoothly after an appropriate delay.
@@
A.2 Energy Cost Function ($C_{2}$)
@@
A.3 Carbon Emission Cost Function ($C_{3}$)
@@
A.4 Immediate and delayed reward responses
@@
Appendix B Proto Definitions
Here, we will elaborate on the exact proto definitions used in the dataset.
@@ Appendix B Proto Definitions
As mentioned above, this data is stored in protos. This section provides the definition of each proto, categorizing them using the four categories above, with examples of each.
B.1 Environment Data Protos
This is the data that provides, once per environment, details about the environment such as number of devices, and zones, etc. There are two proto definitions:
@@ B.1 Environment Data Protos
    ZoneInfo: The ZoneInfo message defines thermal spaces or zones in the building and provides zone-to-device association, which enables using the associated VAVs’ zone air temperatures to estimate the zone’s temperature.
@@
    DeviceInfo: The HVAC devices in the building are defined in the DeviceInfo message. Each device exposes a map of observable_fields and action_fields. The observable_fields represent the observable state of the building in native units, and the action_fields are available setpoints exposed by the building that the agent may add to its action space. Currently observable_fields and action_fields are floating point values, but may be expanded to categorical values in the future.
B.1.1 ZoneInfo Definition
1message ZoneInfo {
2
3
4 enum ZoneType {
5 UNDEFINED = 0;
6 ROOM = 1;
7 FLOOR = 2;
8 OTHER = 10;
9 }
10 // Unique Identifier of the zone.
11 string zone_id = 1;
12 // ID of the building
13 string building_id = 2;
14 // Free-form description of the zone, like microkitchen, office, etc.
15 string zone_description = 3;
16 // Square footage of the zone.
17 float area = 4; // square meters
18 // Zero to multiple device identifiers associated with this zone, like VAVs.
19 repeated string devices = 5;
20 // Optional field to describe the type of zone.
21 ZoneType zone_type = 6;
22 // Optional field to indicate the floor of the building.
23 int32 floor = 7;
24}
B.1.2 ZoneInfo Example
1zone_id: "rooms/9028552253"
2building_id: "buildings/3616672508"
3zone_description: "US-BLDG-2-C201"
4devices: "2614466029028994"
5devices: "2687242320524339"
6devices: "2640423556868160"
7zone_type: ROOM
8floor: 2
B.1.3 DeviceInfo Definition
1// Details about a specific device in the building.
2message DeviceInfo {
3 // Device types in smart buildings (official Carson top-level device types).
4 enum DeviceType {
5 UNDEFINED = 0;
6 FAN = 1;
7 PMP = 2;
8 FCU = 3;
9 VAV = 4;
10 DH = 5;
11 AHU = 6;
12 BLR = 7;
13 CDWS = 8;
14 CH = 9;
15 CHWS = 10;
16 CT = 11;
17 DC = 12;
18 DFR = 13;
19 DMP = 14;
20 HWS = 15;
21 HX = 16;
22 MAU = 17;
23 SDC = 18;
24 UH = 19;
25 PWR = 20;
26 GAS = 21;
27 AC = 22;
28 OTHER = 23;
29 }
30
31
32 enum ValueType {
33 VALUE_TYPE_UNDEFINED = 0;
34 VALUE_CONTINUOUS = 1;
35 VALUE_INTEGER = 2;
36 VALUE_CATEGORICAL = 3;
37 VALUE_BINARY = 4;
38 }
39
40
41 // Unique device identifier.
42 string device_id = 1;
43 // If applicable, the zone associated with the device (like VAVs).
44 string namespace = 2;
45 string code = 3;
46 string zone_id = 4;
47
48
49 // The type of device, VAV, AHU, etc.
50 DeviceType device_type = 5;
51 // Map of measurement name exposed by the device to the value type.
52 map<string, ValueType> observable_fields = 6;
53 // Map of setpoint name exposed by the device to their value type.
54 map<string, ValueType> action_fields = 7;
55}
B.1.4 DeviceInfo Example
1device_id: "202194278473007104"
2namespace: "PHRED"
3code: "US-BLDG:AHU:AC-2"
4device_type: AHU
5observable_fields {
6 key: "building_air_static_pressure_sensor"
7 value: VALUE_CONTINUOUS
8}
9observable_fields {
10 key: "building_air_static_pressure_setpoint"
11 value: VALUE_CONTINUOUS
12}
13action_fields {
14 key: "building_air_static_pressure_setpoint"
15 value: VALUE_CONTINUOUS
16}
17action_fields {
18 key: "cooling_percentage_command"
19 value: VALUE_CONTINUOUS
20}
21action_fields {
22 key: "exhaust_air_damper_percentage_command"
23 value: VALUE_CONTINUOUS
24}
B.2 Observation Data Protos
This includes the measurements from all devices in the building (VAV’s zone air temperature, gas meter’s flow rate, etc.), provided at each time step. There are two proto definitions:
  1. 1.
    ObservationRequest
  2. 2.
    ObservationResponse
To acquire the latest building state, at each timestep the building accepts an ObservationRequest and returns an ObservationResponse. The ObservationRequest contains a UTC timestamp of the requested observation, and list of SingleObservationRequests. Each SingleObservationRequest is a tuple of the device_id and the measurement_name that must match with a device and an observable_field in one of the DeviceInfos exposed by the building. The building returns an ObservationResponse that contains the UTC timestamp from the building, the original ObservationRequest, and a list of SingleObservationResponses. Each SingleObservationResponse contains the associated SingleObservationRequest, the validity time of the measurement/observation, a boolean validity indicator, and the observation, in native units, as a continuous, integer, categorical, binary or string value.
B.2.1 ObservationRequest Definition
1// Agent's request to get the current observation vector.
2message ObservationRequest {
3 // UTC timestamp when the agent generated the request.
4 google.protobuf.Timestamp timestamp = 1;
5 // One or more individual requests.
6 repeated SingleObservationRequest single_observation_requests = 2;
7}
8
9
10// A request to get a single measurement from a specific sensor.
11message SingleObservationRequest {
12 // Unique device identifier.
13 string device_id = 1;
14 // Name of the sensor, e.g., zone_air_temperature.
15 string measurement_name = 2;
16}
B.2.2 ObservationRequest Example
1timestamp {
2 seconds: 1682649309
3 nanos: 942662000
4}
5single_observation_requests {
6 device_id: "202194278473007104"
7 measurement_name: "supply_fan_speed_frequency_sensor"
8}
9single_observation_requests {
10 device_id: "202194278473007104"
11 measurement_name: "mixed_air_temperature_sensor"
12}
13single_observation_requests {
14 device_id: "202194278473007104"
15 measurement_name: "outside_air_flowrate_setpoint"
16}
17single_observation_requests {
18 device_id: "202194278473007104"
19 measurement_name: "supply_air_temperature_sensor"
20}
B.2.3 ObservationResponse Definition
1// Building's response to an observation request message.
2message ObservationResponse {
3 google.protobuf.Timestamp timestamp = 1;
4 ObservationRequest request = 2;
5 repeated SingleObservationResponse single_observation_responses = 3;
6}
7
8
9
10
11// Response for a single observation request.
12message SingleObservationResponse {
13 // The validity time in UTC of the measurement.
14 google.protobuf.Timestamp timestamp = 1;
15 // Original request.
16 SingleObservationRequest single_observation_request = 2;
17 // Validity flag on the observation.
18 bool observation_valid = 3;
19 // Actual observed/measured value.
20 oneof observation_value {
21 float continuous_value = 4;
22 int32 integer_value = 5;
23 string categorical_value = 6;
24 bool binary_value = 7;
25 string string_value = 8;
26 }
27}
B.2.4 ObservationResponse Example
1timestamp {
2 seconds: 1681110000
3}
4request {
5 timestamp {
6 seconds: 1682649309
7 nanos: 942662000
8 }
9 single_observation_requests {
10 device_id: "202194278473007104"
11 measurement_name: "supply_fan_speed_frequency_sensor"
12 }
13 single_observation_requests {
14 device_id: "202194278473007104"
15 measurement_name: "mixed_air_temperature_sensor"
16 }
17 single_observation_requests {
18 device_id: "202194278473007104"
19 measurement_name: "outside_air_flowrate_setpoint"
20 }
21single_observation_responses {
22 timestamp {
23 seconds: 1681109783
24 nanos: 299000000
25 }
26 single_observation_request {
27 device_id: "202194278473007104"
28 measurement_name: "supply_fan_speed_frequency_sensor"
29 }
30 observation_valid: true
31 continuous_value: 0.0
32}
33single_observation_responses {
34 timestamp {
35 seconds: 1681109783
36 nanos: 299000000
37 }
38 single_observation_request {
39 device_id: "202194278473007104"
40 measurement_name: "mixed_air_temperature_sensor"
41 }
42 observation_valid: true
43 continuous_value: 290.3909912109375
44}
45single_observation_responses {
46 timestamp {
47 seconds: 1681109783
48 nanos: 299000000
49 }
50 single_observation_request {
51 device_id: "202194278473007104"
52 measurement_name: "outside_air_flowrate_setpoint"
53 }
54 observation_valid: true
55 continuous_value: 8.825417518615723
56}
B.3 Action Data Protos
This consists of the device setpoint values that the agent wants to set, provided at each timestep. There are two relevant protos:
  1. 1.
    ActionRequest
  2. 2.
    ActionResponse
The Environment converts the action from the agent into an ActionRequest and sends it to the building. The building applies the request and returns an ActionResponse. The ActionRequest contains the UTC timestamp from the Environment, and a list of SingleActionRequests, one for each setpoint in the agent’s action space. Each SingleActionRequest contains a tuple of the device_id, setpoint_name, and requested setpoint_value, in native units. The device_id must match with one of the device_ids in the DeviceInfos, and the setpoint_name must match with one of the action_fields of the associated device. The ActionResponse contains the building’s UTC timestamp, the original ActionRequest, and a list of SingleActionResponses, one associated with each SingleActionRequest. The SingleActionResponse contains the associated SingleActionRequest, a response type enumeration, and a string for additional information.
B.3.1 ActionRequest Definition
1// Agent's request to the building with an action.
2message ActionRequest {
3 // The UTC timestamp that the agent initiated the request.
4 google.protobuf.Timestamp timestamp = 1;
5 // One or more action requests to be performed.
6 repeated SingleActionRequest single_action_requests = 2;
7}
8
9// An action request to assign a value to one setpoint on one device.
10message SingleActionRequest {
11 // The device being commanded.
12 string device_id = 1;
13 // Actual setpoint to be changed, like zone_air_temperature_setpoint.
14 string setpoint_name = 2;
15 oneof setpoint_value {
16 float continuous_value = 3;
17 int32 integer_value = 4;
18 string categorical_value = 5;
19 bool binary_value = 6;
20 string string_value = 7;
21 }
22}
B.3.2 ActionRequest Example
1timestamp {
2 seconds: 1682649309
3 nanos: 942662000
4}
5single_action_requests {
6 device_id: "12945159110931775488"
7 setpoint_name: "supply_air_static_pressure_setpoint"
8 continuous_value: 186.8100128173828
9}
10single_action_requests {
11 device_id: "12945159110931775488"
12 setpoint_name: "supply_air_temperature_setpoint"
13 continuous_value: 294.2592468261719
14}
15single_action_requests {
16 device_id: "13761436543392677888"
17 setpoint_name: "supply_water_temperature_setpoint"
18 continuous_value: 310.9259338378906
19}
20single_action_requests {
21 device_id: "13761436543392677888"
22 setpoint_name: "differential_pressure_setpoint"
23 continuous_value: 82737.09375
24}
25single_action_requests {
26 device_id: "12945159110931775488"
27 setpoint_name: "supervisor_run_command"
28 continuous_value: -1.0
29}
30single_action_requests {
31 device_id: "14409954889734029312"
32 setpoint_name: "supervisor_run_command"
33 continuous_value: -1.0
34}
B.3.3 ActionResponse Definition
1// Building's response to an action request.
2message ActionResponse {
3 // UTC timestamp of the building's response.
4 google.protobuf.Timestamp timestamp = 1;
5 // Original action request.
6 ActionRequest request = 2;
7 // Individual responses for each action.
8 repeated SingleActionResponse single_action_responses = 3;
9}
10
11
12
13// Building's response to a single action request.
14message SingleActionResponse {
15 enum ActionResponseType {
16 UNDEFINED = 0;
17 // The building accepted the action as requested.
18 ACCEPTED = 1;
19 // The building is processing the request, but has not completed.
20 PENDING = 2;
21 // The action request timed out by request handler.
22 TIMED_OUT = 3;
23 // Request is rejected because the set value is not in an acceptable range.
24 REJECTED_INVALID_SETTING = 4;
25 // Rejected because the setting is not enabled or available for control.
26 REJECTED_NOT_ENABLED_OR_AVAILABLE = 5;
27 // A technician or control function overrode the action.
28 REJECTED_OVERRIDE = 6;
29 // The action was assigned to a device that does not exist.
30 REJECTED_INVALID_DEVICE = 7;
31 // The action was assigned to a valid device that's offline.
32 REJECTED_DEVICE_OFFLINE = 8;
33 UNKNOWN = 9;
34 OTHER = 10;
35 }
36
37
38 SingleActionRequest request = 1;
39 ActionResponseType response_type = 2;
40 // Additional optional information related to the action/response.
41 string additional_info = 3;
42}
B.3.4 ActionResponse Example
1timestamp {
2 seconds: 1681110000
3}
4request {
5 timestamp {
6 seconds: 1682649309
7 nanos: 942662000
8 }
9 single_action_requests {
10 device_id: "12945159110931775488"
11 setpoint_name: "supply_air_static_pressure_setpoint"
12 continuous_value: 186.8100128173828
13 }
14 single_action_requests {
15 device_id: "12945159110931775488"
16 setpoint_name: "supply_air_temperature_setpoint"
17 continuous_value: 294.2592468261719
18 }
19 single_action_requests {
20 device_id: "13761436543392677888"
21 setpoint_name: "supply_water_temperature_setpoint"
22 continuous_value: 310.9259338378906
23 }}
24single_action_responses {
25 request {
26 device_id: "12945159110931775488"
27 setpoint_name: "supply_air_static_pressure_setpoint"
28 continuous_value: 186.8100128173828
29 }
30 response_type: ACCEPTED
31 additional_info: "2023-04-10 06:56:23.299000+00:00 12945159110931775488"
32}
33single_action_responses {
34 request {
35 device_id: "12945159110931775488"
36 setpoint_name: "supply_air_temperature_setpoint"
37 continuous_value: 294.2592468261719
38 }
39 response_type: ACCEPTED
40 additional_info: "2023-04-10 06:56:23.299000+00:00 12945159110931775488"
41}
42single_action_responses {
43 request {
44 device_id: "13761436543392677888"
45 setpoint_name: "supply_water_temperature_setpoint"
46 continuous_value: 310.9259338378906
47 }
48 response_type: ACCEPTED
49 additional_info: "2023-04-10 06:55:33.394000+00:00 13761436543392677888"
50}
B.4 Reward Data Protos
This includes information used to calculate the reward, as expressed in cost in dollars, carbon footprint, and comfort level of occupants, provided at each time step The Reward protos define the input and output messages for our 3C reward function (Cost Carbon and Comfort), which contains the code that converts them into a single scalar value, a requirement for most RL algorithms. There are two relevant protos:

### v2

@@ 7 Learning an Improved Control Policy
Here we provide an example of training a Reinforcement Learning Agent on the simulator to generate an improved policy over the current rule-based baseline programmed by the building operators, using building SB1. Inspired by prior effective use of Soft Actor Critic (SAC)(Haarnoja et al., 2018) on related problems (Kathirgamanathan et al., 2021; Coraci et al., 2021; Campos et al., 2022; Biemann et al., 2021), we chose to demo our benchmark using a SAC agent, and compared our agent with the baseline performance of running the policy currently used in the real building. Both actor and critic were feedforward networks. We ran hyperparameter tuning, again using the method from Golovin et. al. (Golovin et al., 2017), to choose the dimensionality of the critic network and actor network, the batch size, the critic learning rate and actor learning rate, and $\gamma$.
@@
We recorded the actor loss, critic loss, alpha loss, and return, over a two-day period. The agents trained for 4,000 iterations. Using the $R_{3C}$ reward, the baseline over this two-day period had a return of -12.9, and our best agent had an improved return of -11.9, an 8% improvement over the baseline, as shown in Table 3. For further details, and a performance comparison between the learned policy and baseline, with a breakdown on setpoint deviation, carbon emissions, electricity, and natural gas, see Appendix J.
@@ 8 Limitations and Conclusion
The biggest limitation of our benchmark is that our simulator lacks a radiative heat model, and we hope further work can add this. In addition, our calibration focused on temperature, but in future work, we hope to include energy consumption metrics as part of the calibration procedure. While our benchmark is grounded in real data, we do not have results on training a model and deploying it on these buildings, something we leave for future work.
@@
We present a high-quality interactive HVAC Control Suite, with an explicit focus on building and benchmarking solutions that scale. Our benchmark has three parts, representing the state of the art of open source HVAC data, scalable data-driven simulation, and physically informed dynamics models. We believe this benchmark will facilitate collaboration, reproducibility, and progress on this problem, making an important contribution towards environmental sustainability.
Impact Statement
This paper has a large potential impact on environmental sustainability. Buildings contribute almost 40% of carbon emissions in the US, with commercial buildings roughly 17%(EIA, ). About half the emissions from commercial buildings come from HVAC(Pérez-Lombard et al., 2008), or roughly 9% of all carbon emissions in the US. Even a tiny reduction will have a massive impact, assuming the solution is one that is able to scale. Our work presents a significant step forward in providing a benchmark to facilitate collaboration, measure progress, and push forward HVAC control optimization with a clear focus on scaling out of the lab and into buildings everywhere.
While this benchmark does not directly enable deployment of policies on real building equipment, such deployment would need to be done with caution and care to mitigate safety concerns and potential damage to hardware. The Smart Buildings simulator, with diverse building settings, helps to address these concerns by allowing the field to study performance across generalization gaps.
Appendix A Learning an Optimal Control Policy
Reinforcement Learning (RL) is a branch of machine learning that attempts to train an agent to choose the best actions to maximize the expected long-term, cumulative reward (Sutton & Barto, 2018). The set of parameters $\boldsymbol{\theta}^{*}$ of the optimal policy can be expressed as:
$\boldsymbol{\theta}^{*}=\arg\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta}(\tau)}\left[\sum_{t}\gamma^{t}R(S_{t},A_{t})\right]$
where $\theta$ is the current policy parameter, and $\tau$ is a trajectory of states, actions, and rewards over sequential timesteps $t$. Over time, the agent explores the action space and learns to maximize the reward over the long term for each given state. A discount factor $\gamma$ reduces the value of future rewards amplifying the value of the near-term reward. When this cycle is repeated over multiple episodes, the agent converges on a state-action policy that maximizes the long-term reward. To converge to the optimal policy, the agent requires many iterations to explore the policy space, making online training directly on the real-world building inefficient, dangerous and impractical. Therefore, it is necessary to enable offline learning, where the agent can train in an efficient sandbox environment that adequately emulates the dynamics of the building before being deployed to the real world.
When applying RL to find an optimal policy in a complex dynamical environment such as a building, there are generally two possible approaches. The model-free RL approach involves learning a policy by directly following gradient signal from the reward function, similar to error backpropagation from a loss function. This is a seemingly straightforward approach but care must be taken to accommodate the stochastic nature of the MDP and the environment, as well as other sources of noise such as sensor noise and delays. A variety of statistical smoothing techniques are often used, generally leading to a slow convergence rate. This process can sometimes get stuck in local optima, oscillate in cycles, or be too slow to keep up with a changing environment.
The alternative model-based RL process involves learning an internal model (or making use of an existing one, such as our simulator or PINN) that predicts the likely state that will result from certain state+action combinations. This internal model (or models) can then be used by an optimizer/learner to explore potential policies in deeper and more sophisticated ways, sometimes even using multi-step look-ahead and other heuristic search schemes suitable for highly rugged reward landscapes or environments with very long-term nonlinear rewards. The learned internal model can be simple, based only on current state and action, or deep, in that it will consider long and short term history and other factors such as weather predictions provided by other models. Model based RL can potentially make better use of past experiences, but in turn requires more computation and could potentially converge prematurely on wrong strategies. Note that both model-based and mode-free RL must allow for some off-policy exploration in order to learn the landscape, leading to the exploration-exploitation dilemma.
Ideally, the internal model used by model-based RL systems – like pre-trained generative systems in general – can contain knowledge that is potentially transferable across tasks. For example, once an internal model can predict the state resulting from certain actions, it can be used for a variety of tasks, such as heating, cooling, minimizing boiler fluctuations, or any arbitrary objective that can be calculated on a trajectory. A dynamics model can also be used to optimize non-RL policies, such as heuristic rule tables, decision trees, or even conventional PID controllers. We view this kind of classical model-predictive control (MPC) as a baseline, as we hope RL agents will learn to outperform them.
Appendix B Reward Function Details
@@
We call our reward function the 3C Reward, because it is made up of a combination of three factors: Comfort, Cost, and Carbon. The purpose of the reward function is to provide the agent with a feedback signal after each action about the quality of the current and past actions performed. We combine the different objectives as a normalized, weighted sum of maintaining comfort conditions, electrical cost, and carbon cost:
@@
where $C_{1}$ represents normalized comfort conditions, $C_{2}$ normalized energy cost and $C_{3}$ normalized carbon emission. Constants $u$, $v$, $w$ represent operator preferences, allowing them to weigh the relative importance of cost, comfort and carbon consumption.
Each value $C_{1},C_{2},C_{3}$, is bounded by the range $[-1,0]$, where worst performance is $-1$ and the ideal performance upper-bound is 00 Thus the reward function in an aggregate is formulated as an approximate regret function, bounded in the range [-1,0], and represents an offset from the best case where comfort conditions are perfectly maintained, without consuming energy and emitting carbon. Each of the sub functions $C_{1},C_{2},C_{3}$ will be elaborated next.
B.1 Comfort Loss Function ($C_{1}$)
@@
Studies have shown that a relationship exists between work performance and temperature and air quality (Deng et al., 2024). For example, in Seppanen et al. (2006), work performance was quantified as the mean time required to complete common office tasks (e.g., text processing, bookkeeping calculations, telephone customer service calls, etc.). Performance was shown to increase gradually with temperatures increasing up to 21-22℃ and decreasing at temperatures beyond 23-24℃. Therefore, when temperatures deviate outside setpoints, the comfort loss should also be smooth and monotonically increasing.
@@
The comfort loss function represents a bounded penalty term for occupied zones that have zone air temperatures outside of setpoint and covers three adjacent temperature intervals: below cooling setpoint $T_{z}<\hat{T}_{heating}$, inside setpoints $\hat{T}_{heating}\leq T_{z}\leq\hat{T}_{cooling}$, and above cooling setpoint $\hat{T}_{cooling}<T_{z}$
@@
Live Occupant Feedback The idea of human feedback shaping the agent’s policy may be particularly suitable for the smart buildings project and has been detailed in Knox and Stone 2009. While not implemented in the initial version of the reward function, the comfort loss function can be extended with an occupant feedback signal reflecting discomfort (e.g., “too hot” or “too cold”) in a variety of methods like Mozer 1998 (Mozer, 1998). The agent’s goal should be to minimize this type of feedback, and the regret should be increased anytime this feedback signal is received. Suppose one or more occupants in zone $z$, provided a “too cold” feedback signal, $\hat{T}_{heating}$ may be increased by a small amount from the baseline setpoint configuration, and may smoothly return to the baseline smoothly after an appropriate delay.
@@
B.2 Energy Cost Function ($C_{2}$)
@@
B.3 Carbon Emission Cost Function ($C_{3}$)
@@
B.4 Immediate and delayed reward responses
@@
B.5 Indoor Air Quality
Indoor air quality is important for multiple reasons, from mitigating exposure to airborne viruses to managing humidity. While the action space in this paper focuses on the hot water and air handler supply temperatures, the amount of fresh air being brought into the air handler is an important factor for air quality. Most air handler units will have a minimum amount of fresh air as specified by a minimum outside air damper position, and the damper position is dependent on the supply air temperature setpoint. Nevertheless, when deploying air handler policies, it is important to verify that the building is receiving sufficient fresh air.
The ANSI/ASHRAE Standard 62.1 standard for indoor air quality (ASHRAE, 2022) recommends the following equation for minimum outside airflow:
$V_{min}=R_{p}\cdot P_{z}+R_{a}\cdot A_{z}$ (3)
Where $R_{p}$ is the rate per person, $P_{z}$ is the zone population, $R_{a}$ is the rate per area, and $A_{z}$ is the zone area. $R_{p}$ for office spaces is 5 CFM per person, but other types of spaces may require higher rates of 10-20 CFM per person, or 30 CFM per person for infection risk management as specified by ASHRAE Standard 241-2023 (Sherman & Jones, 2023). The $R_{a}$ for office spaces is 0.06 CFM per square foot, but this could similarly be increased for different space types.
There are two ways to incorporate this standard: one is by adding a reward term and the other is by applying a constraint at deployment time, both described below.
Modified Reward and Action Space An air quality term could be added to the 3C reward to penalize states with lower outside airflow. Similar to the comfort reward term, this could use a logistic sigmoid to represent a smooth, bounded decay, with 0.5 loss being the point where outside airflow $V_{OA}$ equals the minimum standard $V_{min}$ as defined in equation 3.
$R(V_{OA})=\frac{1}{1+e^{-\lambda(V_{OA}-V_{min})}}$ (4)
Parameter $\lambda$ is a stiffness coefficient that affects the slope of decay, and might have a value between 1 and 10 depending on the requirements of the deployment space.
To better optimize the overall system, the static pressure setpoint and outside air damper position could be added to the action space. Another way to incorporate air quality would be to add a feasibility constraint instead of a reward term during policy learning (Chen et al., 2021).
Constraining at Deployment When deploying a learned 3C policy, one can calculate the minimum outside airflow constraint (equation 3) using the deployment zone’s occupancy and square footage. The outside air damper position can be set to increase when the outside airflow, as measured by static pressure sensors, drops below this constraint. Another way to incorporate indoor air quality would be to combine it with predictive control (Wang & Dong, 2023).
Appendix C Dataset Format
C.1 Benchmark Access
The data is part of Tensorflow Datasets (TFDS), and can be downloaded from [link redacted for blind review].
The code is available at [link redacted for blind review].
C.2 Dataset Format
Here, we will elaborate on the exact format of the dataset.
@@ C.2 Dataset Format
C.3 Environment Data
This is the data that provides, once per environment, details about the environment such as number of devices or zones. There are four types of data: ZoneInfo, DeviceInfo, Floorplan, and Device Locations.
@@ C.3 Environment Data
    ZoneInfo: The ZoneInfo defines thermal spaces or zones in the building and provides zone-to-device association, which enables using the associated VAVs’ zone air temperatures to estimate the zone’s temperature.
@@
    DeviceInfo: The HVAC devices in the building are defined in the DeviceInfo. Each device exposes a map of observable_fields and action_fields. The observable_fields represent the observable state of the building in native units, and the action_fields are available setpoints exposed by the building that the agent may add to its action space. Currently observable_fields and action_fields are floating point values, but may be expanded to categorical values in the future.
  3. 3.
    Floorplan: This is a 2d matrix, where there are 4 possible values in each cell: outside air, inside air, exterior wall, interior wall.
  4. 4.
    Device Locations: This is a dictionary of device names to cells in the floorplan. For each device, the map provides the cells corresponding to the room that the device is located in.
C.4 Observation Data
This includes the measurements from all devices in the building (VAV’s zone air temperature, gas meter’s flow rate, etc.), provided at each timestep. This data is given as a matrix of TxM, where there are T timesteps and M measurements per timestep. A list of T timestamps is provided, as well as a list of M measurement names, defining the rows and columns of the matrix.
C.5 Action Data
This consists of the device setpoint values that the agent wants to set, provided at each timestep. This data is given as a matrix of TxS, where there are T timesteps and S setpoints per timestep. A list of T timestamps is provided, as well as a list of S setpoint names, defining the rows and columns of the matrix.
C.6 Reward Data
This includes information used to calculate the reward, as expressed in cost in dollars, carbon footprint, and comfort level of occupants, provided at each timestep The reward data is further divided into two categories: